OpenStudy (anonymous):

The formula for the amount with simple interest is A = P(1 + rt). Find r, if A = 224, P = 200, and t = 2.

OpenStudy (anonymous):

Since you are given the formula, all you need to do is substitute the numbers into the equation and solve for r. A=P(1+rt) 224=200(1+r(2)) 224=200+400r 24=400r r=.06 Then you can check to see if that is the right answer by plugging it into the equation A=200(1+.06*2) A=224